Problem 5.19 (Excel Video) You decide to take advantage of the current online dating craze and start your own Web site. You know that you have 400 people who will sign up immediately and, through a careful marketing research and analysis, determine that membership can grow by 29 percent in the first two years, 20 percent in year 3, and 18 percent in Year 4. How many members do you expect to have at the end of four years? (Round intermediate calculations to 6 decimal places, in all cases round your final answer to the nearest whole number) Excel Template (Note: This template includes the problem statement as it appears in your textbook. The problem assigned to you here may have different values. When using this template, copy the problem statement from this screen for easy reference to the values youve been given here, and be sure to update any values that may have been pre-entered in the template based on the textbook version of the problem.) Members at the end of four years
Answer #1
Members at end of 4 Years                             943
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
Member at end of year 1 = 400 * 1.29              516.000000
Member at end of year 2 = 516*1.29              665.640000
Member at end of year 3 = 665.64*1.20              798.768000
Member at end of year 4 = 798.768*1.18              942.546240